AN OVERVIEW OF PNC RELATED STUDIES INVOLVING HEAVY POLAR MOLECULES

\underline{TIMOTHY C. STEIMLE} , Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

Parity non-conservation studies (PNC) has given added impetus for both experimental and theoretical determination of the properties of heavy polar molecules. An overview of the current experimental approaches used by various groups in an attempt to the determine the electron electric dipole moment (eEDM), d_{e}, and the anapole moment of nuclei will presented. The status of the characterization of the various molecules proposed as venues in PNC will be summarized.